OpenStudy (anonymous):

How would I go about doing this? Solve each quadratic equation. Show your work. 14. (2x – 1)(x + 7) = 0

OpenStudy (anonymous):

You can use the zero product property: If two things have a product of zero, then at least one of them must be zero. Symbolically: ab=0 implies a=0 or b=0

OpenStudy (anonymous):

I'm assuming the 14. is the problem number and not part of the problem itself: If you have (2x-1)(x+7)=0, then a is 2x-1 and b is x+7. This means that 2x-1=0 or x+7=0. This will give you two solutions - one for each equation.

OpenStudy (anonymous):

Interesting.

OpenStudy (anonymous):

If you had something more complicated, say \[(5x+9)(2x-3)(4x)(3x-2)=0\]Then you would make each set of parentheses (four in this case) equal to zero and solve them individually. It's a clever use of factoring and the zero product property.

Directrix (directrix):

As stated above: (2x – 1) (x + 7) = 0 which means 2x - 1 = 0 or x + 7 = 0 Solve each equation separately for x. @Jonnychewy Post what you get and somebody will check it for you, okay?

OpenStudy (anonymous):

Okie dokie! Thank you for explaining!

OpenStudy (anonymous):

No worries :)

Directrix (directrix):

Okay. The idea is (as the other poster stated) that if two quantities multiply to zero, at least one of them has to be zero. So, in this case, it is 2x-1 or x + 7 that is 0. Your task is the find the value of x that makes each one of those factors equal to 0. Go ahead and solve them now and just be done with this problem.

OpenStudy (anonymous):

Do you think you could give me some steps, @directrix ?

OpenStudy (anonymous):

Or @iGreen ...

OpenStudy (anonymous):

Could you help me perhaps @iGreen ?

OpenStudy (anonymous):

1/2 and -7 ?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (turingtest):

that's correct

OpenStudy (shaik0124):

u can get a quadratic equation too